abbidoo:

the sum of 2 consecutive integers is 226 what is the largest consecutive integer?

Hoodmemes:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @dontsaymyname Heyy!! so basically we represent the first integer with y and the second even integer equal to y + 2 ↓ Thus using this information we can complete a system of equations ↓ x + (x + 2) = 226 Combine like terms 2x + 2 = 226 Subtract 2 from both sides 2x = 224 Divide x = 112 The two integers are 112 and 114. You can double check by them 112 + 114 = 226 Hopefully this helps! Feel free to click the Best Response button by my answer, I would greatly appreciate it :) \(\color{#0cbb34}{\text{End of Quote}}\) wow u did all that?

abbidoo:

thank you!

jhonyy9:

2 consecutive integers - what mean for you ?

abbidoo:

what do you mean?

jhonyy9:

what are two consecutive integers for example ?

jhonyy9:

@Vocaloid

Hoodmemes:

Didnt she just solve the problem for the user?

jhonyy9:

but this not is correct solved

Hoodmemes:

oo

jhonyy9:

two consecutive integers are 2 and 3 for example one odd and one even

jhonyy9:

@abbidoo please repost this problem and tag me there so i will help you solve it correct

jhonyy9:

@Shadow opinion please

jhonyy9:

ty

dontsaymyname:

@Laylalyssa

Shadow:

Yeah it should be x + (x+1) = no? A consecutive number is just +1 or -1 depending how you look at it. The problem is that when you get 2x + 1 = 226, simplified to 2x = 225, x = 225/2, which is not an integer. When I googled the question, it came up with 'even consecutive integers,' so I think the poster forgot that part. It would make x + (x + 2) make sense.

Shadow:

1 attachment
Shadow:

Posted a screenshot of the google results

jhonyy9:

ok i understand it now - just the text of the problem not is correct

jhonyy9:

ty